# OctFusion: Octree-based Diffusion Models for 3D Shape Generation

### v1

@@ Abstract.
submissionid: 293journal: TOGjournalyear: 2024copyright: acmcopyrightccs: Computing methodologies Shape modelingccs: Computing methodologies Neural networks
Figure 1. OctFusion is capable of generating high-quality and high-resolution 3D shapes in various scenarios, such as unconditional/label-conditional generation, text/sketch-guided generation, and textured mesh generation.
1. Introduction
3D content creation is a fundamental task in computer graphics and has a broad range of applications, such as virtual reality, augmented reality, 3D games, and movies. Recently, generative neural networks, especially diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), have achieved remarkable progress in 3D generation and have attracted much attention in academia and industry.
However, it is still challenging for diffusion models to efficiently generate highly detailed 3D shapes. Several works seek to generate 3D shapes by distilling multiview information contained in well-trained 2D diffusion models (Poole et al., 2023; Lin et al., 2023), which involves a costly per-shape optimization process and requires minutes, even hours, to generate one single 3D output. Moreover, the generated results often suffer from the multi-face Janus problem and may contain oversaturated colors. On the other hand, many works (Zheng et al., 2023; Cheng et al., 2023; Hui et al., 2022; Li et al., 2023; Chou et al., 2023; Zhang et al., 2023; Gupta et al., 2023) directly train diffusion models on 3D shape datasets. Although these methods can generate 3D shapes in several seconds by directly forwarding the trained 3D models, the results are often of low resolution due to the limited expressiveness of shape representations like Triplanes (Gupta et al., 2023; Shue et al., 2023) or high computational and memory cost of 3D neural networks. To generate fine-grained geometric details, multi-stage latent diffusion models (Rombach et al., 2022; Cheng et al., 2023; Zhang et al., 2023) or cascaded training schemes (Zheng et al., 2023; Ren et al., 2024) have been introduced, which further increases the complexity of the training process.
@@ 1. Introduction
For the shape representation, we represent a 3D shape as a volumetric octree and append a latent feature on each leaf node. The latent features are decoded into local signed distance fields (SDFs) with a shared MLP, which are then fused into a global SDF by multi-level partition-of-unity (MPU) modules (Wang et al., 2022; Ohtake et al., 2003). This representation combines the benefits of both implicit representations (Chen and Zhang, 2019; Park et al., 2019; Mescheder et al., 2019) for representing continuous fields and explicit spatial octree structures (Wang et al., 2022) for expressing complex geometric and texture details. The octree can be constructed from a point cloud or a mesh by recursive subdivision; the latent features are extracted with a variational autoencoder (VAE) (Kingma and Welling, 2013) built upon dual octree graph networks (Wang et al., 2022). The representation can also be extended to support color fields by additional latent features. The SDF and color fields can be converted to triangle meshes paired with an RGB color on each vertex as output with the marching cube algorithm (Lorensen and Cline, 1987).
@@
To train diffusion models on the octree-based representation, our key insight is to regard the splitting status of octree nodes as 0/1 signals; then, we add noise to both the splitting signal and the latent feature defined on each octree node to get a noised octree. The diffusion model essentially trains a U-Net (Ronneberger et al., 2015) to revert the noising process to predict clean octrees from noised octrees for generation. Since the noise is added to all octree levels, a natural idea is to adopt cascaded training schemes (Zheng et al., 2023; Ren et al., 2024) to train a separate diffusion model on each octree level to predict the splitting signals and the final latent features in a coarse-to-fine manner. However, training multiple diffusion models is complex and inefficient, especially for deep octrees. Our key observation is that the octree itself is hierarchical; when generating the deep octree nodes, the shallow nodes have already been generated, resulting in nested U-Net structures for different octree levels. Based on this observation, we propose to train a unified diffusion model for different octree levels, which reuses the trained weights for shallow octree levels nodes when denoising deep octree nodes. Our diffusion model enables weights and computation sharing across different octree levels, thus significantly reducing the parameter number and the training complexity and making our model capable of generating detailed shapes efficiently.
@@
We verify the efficiency, effectiveness, and generalization ability of OctFusion on the widely-used ShapeNet dataset (Chang et al., 2015). OctFusion achieves state-of-the-art performances with only 33M trainable parameters. The generated implicit fields are guaranteed to be continuous and can be converted to meshes with arbitrary resolutions. OctFusion can predict a mesh in less than 2.5 seconds on a single Nvidia 4090 GPU under the setting of 50 diffusion sampling steps. We also train OctFusion on a subset of Objverse (Deitke et al., 2023) and verify it has a strong ability to generate shapes from a complex distribution. We further extend OctFusion to support conditioned generation from sketch images, text prompts, and category labels, where OctFusion also achieves superior performances compared with previous methods (Zheng et al., 2023; Cheng et al., 2023). In summary, our main contributions are as follows:
@@
    We present an octree-based latent representation for 3D textured shapes, which combines the benefits of both implicit representations and explicit spatial octree structures.
@@
2. Related Work
3D Shape Representations
@@ 2. Related Work
Different from images that are often defined on regular grids, 3D shapes have different representations. Early works (Wu et al., 2015, 2016; Zheng et al., 2022) represent 3D shapes as uniformly sampled voxel grids, with which image generative models can be directly extended to the 3D domain. However, voxel grids incur huge computational and memory costs; thus, these methods can only generate 3D shapes with low resolutions. To improve the efficiency, sparse-voxel-based representations, like octrees (Wang et al., 2018, 2022) or Hash tables (Choy et al., 2019; Müller et al., 2022) are proposed to represent 3D shapes with only non-empty voxels, which enables the generation of high-resolution 3D shapes. Another type of 3D representation is point clouds. Due to the flexibility and efficiency of point clouds, they are widely used for 3D generation (Fan et al., 2017; Luo and Hu, 2021; Liu et al., 2019; Nichol et al., 2022; Zeng et al., 2022). However, point clouds are discrete and unorganized; thus, additional efforts are needed to convert point clouds to continuous surfaces, which are more desirable for many graphics applications. To model continuous surfaces of 3D shapes, MLP-based distance or occupancy fields are proposed as implicit representations of 3D shapes (Park et al., 2019; Chen and Zhang, 2019; Mescheder et al., 2019). Although these methods can represent 3D shapes with infinite resolutions, they are computationally expensive since each query of the fields requires a forward pass of the MLP. Recently, triplanes (Peng et al., 2020; Shue et al., 2023) are combined with MLPs to further increase the expressiveness and efficiency, whereas it is observed that triplanes are still hard to model complex geometric details (Zhang et al., 2023; Wang et al., 2022; Zhang et al., 2022). Our shape representation extends the neural MPU in (Wang et al., 2022) and combines the benefits of both implicit representations and octrees, which can represent continuous fields and model complex geometric and texture details efficiently.
3D Diffusion Models
@@
Recently, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have demonstrated great potential for generating diverse and high-quality samples in the image domain (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022), surpassing the performance of GANs (Goodfellow et al., 2016; Dhariwal and Nichol, 2021) and VAEs (Kingma and Welling, 2013). Following this progress, a natural idea is to extend them to the 3D domain. 3D diffusion models with point clouds (Nichol et al., 2022; Luo and Hu, 2021; Zhou et al., 2021) or voxel grids (Li et al., 2023; Hui et al., 2022; Chou et al., 2023; Shim et al., 2023) are first proposed and have achieved promising shape generation results. However, the efficiency and quality of generated shapes are relatively low. Inspired by latent diffusion models (Rombach et al., 2022), many follow-up works also train diffusion models on the latent space of 3D shapes, and the latent space is often obtained with a VAE trained on voxels (Cheng et al., 2023), point clouds (Zeng et al., 2022), triplanes (Gupta et al., 2023; Shue et al., 2023), or implicit shape representations (Zhang et al., 2023; Jun and Nichol, 2023; Erkoç et al., 2023). Another strategy for efficiency and quality improvement is to leverage sparse-voxel-based representations, like octree (Zheng et al., 2023), and adopt cascaded training schemes to generate sparse voxels at different resolutions (Zheng et al., 2023). Subsequent works continue to train cascaded models on the larger dataset (Liu et al., 2023b) or increase the number of training stages for higher resolutions (Ren et al., 2024). Different from these methods, our OctFusion is trained with an unified U-net and can generate 3D shapes with $1024^{3}$ resolutions in a single network, which significantly reduces the complexity of the training procedure.
3D Generation with 2D Diffusion Priors
@@
Apart from training 3D diffusion models, another line of research attempts to distill multi-view image priors from 2D diffusion models to generate 3D shapes, popularized by DreamFusion (Poole et al., 2023) and Magic3D (Lin et al., 2023). The key idea is to optimize 3D representations such as NeRF (Mildenhall et al., 2020) or InstantNGP (Müller et al., 2022) with the gradient guidance from 2D diffusion priors (Poole et al., 2023; Wang et al., 2023). Although these methods can generate diverse and high-quality 3D shapes without access to 3D training data, they are computationally expensive since the optimization of 3D representations often takes minutes or even hours; and the resulting textures are often oversaturated, and the generated shapes contain artifacts due to the inconsistency of multiview image priors provided by 2D diffusion models. Many follow-up works (Chen et al., 2023; Qian et al., 2023; Deng et al., 2023; Tang et al., 2023; Shi et al., 2023) further improve the efficiency and 3D consistency. Our OctFusion can generate 3D shapes in a feed-forward manner within several seconds.
Unified Diffusion Models
@@
To enable diffusion models to generate high-resolution images, a commonly-used strategy is to train cascaded diffusion models (Ho et al., 2022; Nichol and Dhariwal, 2021; Saharia et al., 2022; Ramesh et al., 2022). To reduce the complexity of the training process and share weights and computation across different levels of resolutions, several approaches (Hoogeboom et al., 2023; Jabri et al., 2022; Chen, 2023; Gu et al., 2023) propose to train a unified diffusion model to directly generate high-resolution images. UniDiffuser (Bao et al., 2023) proposes a unified diffusion model to model the joint distribution of multi-modal data, which can generate diverse and high-quality samples with different modalities. The design of OctFusion is inspired by these pioneering works which training an unified U-Net for different resolutions.
3. Method
@@
Figure 2. Overview. Given a point cloud, possibly with color information, an octree is firstly constructed. Then, a VAE is trained to learn latent features on all octree leaf nodes, which can be decoded to continuous distance fields and color fields. Next, the first stage model $\mathcal{F}_{1}$ is trained to predict splitting signals from its noised version, presented as the noised and predicted octree in the figure. Image or text conditions can be optionally provided to guide the generation. Then second stage model $\mathcal{F}_{2}$ predicts latent features, presented as the colors on octree leaf nodes. Notably, the weights of U-Net are shared across different octree levels, which plays an important role in saving computation and memory as well as improving performance.
Overview
@@
3.1. Octree-based Latent Representation
Shape Representation
We encode 3D shapes with octree-based latent representations, which can be decoded into continuous fields, such as signed distance fields (SDFs). Given a mesh or a point cloud, we convert it to an octree by recursive subdividing nonempty octree nodes until the maximum depth is reached. All leaf nodes of the octree form an adaptive partition of the 3D volume. Inspired by (Wang et al., 2022; Ohtake et al., 2003), we store latent features on each leaf node, which are decoded to local SDFs by a shared MLP. Then, we blend local SDFs into global continuous fields using a multi-level partition-of-unity (MPU) module. A 2D illustration of the proposed representation is shown in Fig. 3.
@@ 3.1. Octree-based Latent Representation
(1) $F_{sdf}(p)$ $=\frac{\sum_{i}w_{i}(x)\cdot\Phi_{sdf}(x,f_{i})}{\sum_{i}w_{i}(x)},$
where $x=(p-o_{i})/r_{i}$, representing the local coordinates of $p$ relative to $o_{i}$, $w_{i}(x)$ is a locally-supported linear B-Spline function, and $\Phi_{sdf}(x,f_{i})$ is a shared MLP that maps the local coordinates $x$ and the latent feature $f_{i}$ to the SDF value at $p$. Since $w_{i}(x)$ and $\Phi_{sdf}(x,f_{i})$ are all continuous functions, $F_{sdf}(x)$ are guaranteed to be continuous (Ohtake et al., 2003; Wang et al., 2022). $\Phi_{sdf}(x,f_{i})$ is fully differentiable, and its evaluation is efficient since the weight function $w_{i}(x)$ is locally supported.
Octree-based Variational Autoencoders
We use the variational autoencoder (VAE) (Kingma and Welling, 2013) built upon dual octree graph networks (Wang et al., 2022) to learn the octree-based latent representation. The encoder of the VAE takes an octree built from a point cloud as input and outputs a latent feature for each leaf node; the decoder reconstructs the continuous SDF from the latent features.

### v2

@@ Abstract
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371.10010352.10010381</concept_id> <concept_desc>Computing methodologies Collision detection</concept_desc> <concept_significance>300</concept_significance> </concept> <concept> <concept_id>10010583.10010588.10010559</concept_id> <concept_desc>Hardware Sensors and actuators</concept_desc> <concept_significance>300</concept_significance> </concept> <concept> <concept_id>10010583.10010584.10010587</concept_id> <concept_desc>Hardware PCB design and layout</concept_desc> <concept_significance>100</concept_significance> </concept> </ccs2012>
[…]
volume: 44issue: 5
1 Introduction
3D content creation is a fundamental task in computer graphics and has a broad range of applications, such as virtual reality, augmented reality, 3D games, and movies. Recently, generative neural networks, especially diffusion models [Dickstein2015, Ho2020], have achieved remarkable progress in 3D generation and have attracted much attention in academia and industry.
However, it is still challenging for diffusion models to efficiently generate highly detailed 3D shapes. Several works seek to generate 3D shapes by distilling multiview information contained in well-trained 2D diffusion models [Poole2022, Lin2023], which involves a costly per-shape optimization process and requires minutes, even hours, to generate one single 3D output. Moreover, the generated results often suffer from the multi-face Janus problem and may contain oversaturated colors. On the other hand, many works [Zheng2023, Cheng2023, Hui2022, Li2023, Chou2023, Zhang2023a, Gupta2023] directly train diffusion models on 3D shape datasets. Although these methods can generate 3D shapes in several seconds by directly forwarding the trained 3D models, the results are often of low resolution due to the limited expressiveness of shape representations like Triplanes [Gupta2023, Shue2023] or high computational and memory cost of 3D neural networks. To generate fine-grained geometric details, multi-stage latent diffusion models [Rombach2022, Cheng2023, Zhang2023a] or cascaded training schemes [Zheng2023, Ren2024] have been introduced, which further increases the complexity of the training process.
@@ 1 Introduction
For the shape representation, we represent a 3D shape as a volumetric octree and append a latent feature on each leaf node. The latent features are decoded into local signed distance fields (SDFs) with a shared MLP, which are then fused into a global SDF by multi-level partition-of-unity (MPU) modules [Wang2022, Ohtake2003]. This representation combines the benefits of both implicit representations [Chen2019, Park2019, Mescheder2019] for representing continuous fields and explicit spatial octree structures [Wang2022] for expressing complex geometric and texture details. The octree can be constructed from a point cloud or a mesh by recursive subdivision; the latent features are extracted with a variational autoencoder (VAE) [Kingma2013] built upon dual octree graph networks [Wang2022]. The representation can also be extended to support color fields by additional latent features. The SDF and color fields can be converted to triangle meshes paired with an RGB color on each vertex as output with the Marching Cubes algorithm [Lorensen1987]. Although some previous methods also utilized octree [Zheng2023] or sparse voxel [Ren2024, xiang2024structured] to achieve high-resolution 3D shape generation, their representation can not guarantee the completeness and continuity of the generated implicit field. On the contrary, our dual octree graph representation forms a complete coverage of the 3D bounding volume, which is further discussed in Section 4.2.1.
@@
To train diffusion models on the octree-based representation, our key insight is to regard the splitting status of octree nodes as 0/1 signals; then, we add noise to both the splitting signal and the latent feature defined on each octree node to get a noised octree. The diffusion model essentially trains a U-Net [Ronneberger2015] to revert the noising process to predict clean octrees from noised octrees for generation. Since the noise is added to all octree levels, a natural idea is to adopt cascaded training schemes [Zheng2023, Ren2024] to train a separate diffusion model on each octree level to predict the splitting signals and the final latent features in a coarse-to-fine manner. However, training multiple diffusion models is complex and inefficient, especially for deep octrees. Our key observation is that the octree itself is hierarchical; when generating the deep octree nodes, the shallow nodes have already been generated, resulting in nested U-Net structures for different octree levels. Based on this observation, we propose to train a unified diffusion model for different octree levels, which reuses the trained weights for shallow octree levels nodes when denoising deep octree nodes. Our diffusion model enables weights and computation sharing across different octree levels, thus significantly reducing the parameter number and the training complexity and making our model capable of generating detailed shapes efficiently.
@@
We verify the efficiency, effectiveness, and generalization ability of OctFusion on the widely-used ShapeNet dataset [Chang2015]. OctFusion achieves state-of-the-art performances with only 33M trainable parameters. The generated implicit fields are guaranteed to be continuous and can be converted to meshes with arbitrary resolutions. OctFusion can predict a mesh in less than 2.5 seconds on a single Nvidia 4090 GPU under the setting of 50 diffusion sampling steps. We also train OctFusion on a subset of Objverse [Deitke2023] and verify it has a strong ability to generate shapes from a complex distribution. We further extend OctFusion to support conditioned generation from sketch images, text prompts, and category labels, where OctFusion also achieves superior performances compared with previous methods [Zheng2023, Cheng2023]. In summary, our main contributions are as follows:
@@
    We present an octree-based latent representation for high resolution 3D shapes modeling, which combines the benefits of both implicit representations and explicit spatial octree structures.
@@
2 Related Work
2.1 3D Shape Representations
@@ 2 Related Work
Different from images that are often defined on regular grids, 3D shapes have different representations. Early works [Wu2015, Wu2016, Zheng2022] represent 3D shapes as uniformly sampled voxel grids, with which image generative models can be directly extended to the 3D domain. However, voxel grids incur huge computational and memory costs; thus, these methods can only generate 3D shapes with low resolutions. To improve the efficiency, sparse-voxel-based representations, like octrees [Wang2018a, Wang2022] or Hash tables [Choy2019, Muller2022] are proposed to represent 3D shapes with only non-empty voxels, which enables the generation of high-resolution 3D shapes. Another type of 3D representation is point clouds. Due to the flexibility and efficiency of point clouds, they are widely used for 3D generation [Fan2017, Lou2021, Liu2019a, Nichol2022, Zeng2022]. However, point clouds are discrete and unorganized; thus, additional efforts are needed to convert point clouds to continuous surfaces, which are more desirable for many graphics applications. To model continuous surfaces of 3D shapes, MLP-based distance or occupancy fields are proposed as implicit representations of 3D shapes [Park2019, Chen2019, Mescheder2019]. Although these methods can represent 3D shapes with infinite resolutions, they are computationally expensive since each query of the fields requires a forward pass of the MLP. Recently, triplanes [Peng2020, Shue2023] are combined with MLPs to further increase the expressiveness and efficiency, whereas it is observed that triplanes are still hard to model complex geometric details [Zhang2023a, Wang2022, Zhang2022a]. Our shape representation extends the neural MPU in [Wang2022] and combines the benefits of both implicit representations and octrees, which can represent continuous fields and model complex geometric and texture details efficiently.
2.2 3D Diffusion Models
@@
Recently, diffusion models [Dickstein2015, Ho2020] have demonstrated great potential for generating diverse and high-quality samples in the image domain [Ramesh2022, Saharia2022, Rombach2022], surpassing the performance of GANs [Goodfellow2016a, Dhariwal2021] and VAEs [Kingma2013]. Following this progress, a natural idea is to extend them to the 3D domain. 3D diffusion models with point clouds [Nichol2022, Lou2021, Zhou2021a] or voxel grids [Li2023, Hui2022, Chou2023, Shim2023] are first proposed and have achieved promising shape generation results. However, the efficiency and quality of generated shapes are relatively low. Inspired by latent diffusion models [Rombach2022], many follow-up works also train diffusion models on the latent space of 3D shapes, and the latent space is often obtained with a VAE trained on voxels [Cheng2023], point clouds [Zeng2022], triplanes [Gupta2023, Shue2023], or implicit shape representations [Zhang2023a, Jun2023, Erkoc2023]. Another strategy for efficiency and quality improvement is to leverage sparse-voxel-based representations, like octree [Zheng2023], and adopt cascaded training schemes to generate sparse voxels at different resolutions [Zheng2023]. Subsequent works continue to train cascaded models on the larger dataset [Liu2023, xiang2024structured] or increase the number of training stages for higher resolutions [Ren2024]. Different from these methods, our OctFusion is trained with an unified U-net and can generate 3D shapes with $1024^{3}$ resolutions in a single network, which significantly reduces the complexity of the training procedure.
2.3 3D Generation with 2D Diffusion Priors
@@
Apart from training 3D diffusion models, another line of research attempts to distill multi-view image priors from 2D diffusion models to generate 3D shapes, popularized by DreamFusion [Poole2022] and Magic3D [Lin2023]. The key idea is to optimize 3D representations such as NeRF [Mildenhall2020] or InstantNGP [Muller2022] with the gradient guidance from 2D diffusion priors [Poole2022, Wang2023a]. Although these methods can generate diverse and high-quality 3D shapes without access to 3D training data, they are computationally expensive since the optimization of 3D representations often takes minutes or even hours; and the resulting textures are often oversaturated, and the generated shapes contain artifacts due to the inconsistency of multiview image priors provided by 2D diffusion models. Many follow-up works [Chen2023, Qian2023, Deng2023, Tang2023, Shi2023] further improve the efficiency and 3D consistency. Our OctFusion can generate 3D shapes in a feed-forward manner within several seconds.
2.4 Unified Diffusion Models
@@
To enable diffusion models to generate high-resolution images, a commonly-used strategy is to train cascaded diffusion models [Ho2022a, Nichol2021, Saharia2022, Ramesh2022]. To reduce the complexity of the training process and share weights and computation across different levels of resolutions, several approaches [Emiel2023, Jabri2022, Chen2023a, Gu2023] propose to train a unified diffusion model to directly generate high-resolution images. UniDiffuser [Bao2022] proposes a unified diffusion model to model the joint distribution of multi-modal data, which can generate diverse and high-quality samples with different modalities. The design of OctFusion is inspired by these pioneering works which training an unified U-Net for different resolutions.
3 Method
@@
Figure 1: Overview. Given a point cloud, possibly with color information, an octree is firstly constructed. Then, a VAE is trained to learn latent features on all octree leaf nodes, which can be decoded to continuous distance fields and color fields. Next, the first stage model $\mathcal{F}_{1}$ is trained to predict splitting signals from its noised version, presented as the noised and predicted octree in the figure. Image or text conditions can be optionally provided to guide the generation. Then second stage model $\mathcal{F}_{2}$ predicts latent features, presented as the colors on octree leaf nodes. Notably, the weights of U-Net are shared across different octree levels, which plays an important role in saving computation and memory as well as improving performance.
Figure 2: Octree-based Latent Representation. Here 2D figures are used for better visualization. Left: An input point cloud sampled from a 3D shape. Middle: The octree structure constructed from the point cloud and latent features for octree leaf nodes produced by the VAE. The latent features are shown as solid dots. Right: The continuous SDF field reconstructed decoded from the latent features. The field value of an arbitrary query point $p$ is computed with the MPU module.
3.1 Overview
@@
3.2 Octree-based Latent Representation
3.2.1 Shape Representation
We encode 3D shapes with octree-based latent representations, which can be decoded into continuous fields, such as signed distance fields (SDFs). Given a mesh or a point cloud, we convert it to an octree by recursive subdividing nonempty octree nodes until the maximum depth is reached. All leaf nodes of the octree form an adaptive partition of the 3D volume. Inspired by [Wang2022, Ohtake2003], we store latent features on each leaf node, which are decoded to local SDFs by a shared MLP. Then, we blend local SDFs into global continuous fields using a multi-level partition-of-unity (MPU) module. A 2D illustration of the proposed representation is shown in Fig. 2.
@@ 3.2.1 Shape Representation
$F_{sdf}(p)$ $=\frac{\sum_{i}w_{i}(x)\cdot\Phi_{sdf}(x,f_{i})}{\sum_{i}w_{i}(x)},$ (1)
where $x=(p-o_{i})/r_{i}$, representing the local coordinates of $p$ relative to $o_{i}$, $w_{i}(x)$ is a locally-supported linear B-Spline function, and $\Phi_{sdf}(x,f_{i})$ is a shared MLP that maps the local coordinates $x$ and the latent feature $f_{i}$ to the SDF value at $p$. Since $w_{i}(x)$ and $\Phi_{sdf}(x,f_{i})$ are all continuous functions, $F_{sdf}(x)$ are guaranteed to be continuous [Ohtake2003, Wang2022]. $\Phi_{sdf}(x,f_{i})$ is fully differentiable, and its evaluation is efficient since the weight function $w_{i}(x)$ is locally supported.
3.2.2 Octree-based Variational Autoencoders
We use the variational autoencoder (VAE) [Kingma2013] built upon dual octree graph networks [Wang2022] to learn the octree-based latent representation. The encoder of the VAE takes an octree built from a point cloud as input and outputs a latent feature for each leaf node; the decoder reconstructs the continuous SDF from the latent features.